Your grandmother needs your help. She has $50,000 to invest. Part of this money is to be invested in noninsured bonds paying 15% annual interest. The rest of this money is to be invested in a government-insured certificate of deposit paying 7% annual interest. She told you that she requires a total $6000 per year in extra income from these investments. How much money should be placed in each investment
Answer by mananth(16946)

Part I 15.00% per annum ------------- Amount invested =x
Certificates 7.00% per annum ------------ Amount invested = y
50000
Interest----- 6000

Part I 15.00% per annum ---x
Certificates 7.00% per annum ---y
Total investment
x + 1 y= 50000 -------------1
Interest on both investments
15.00% x + 7.00% y= 6000
Multiply by 2
15 x + 7 y= 600000.00
Multiply (1) by -15
we get
-15 x -15 y= -750000.00
Add this to (2)
0 x -8 y= -150000
divide by -8
y = 18750
Part I 15.00% $ 31250
Certificates 7.00% $ 18750

CHECK
31250 --------- 15.00% ------- 4687.50
18750 ------------- 7.00% ------- 1312.50
Total -------------------- 6000.00

m.ananth@hotmail.ca